OpenStudy (anonymous):

what value of m makes the equation true? 10m + 32 = 48

OpenStudy (anonymous):

10m +32 = 48 subtract 32 from both sides to get: 10m = 16 then divide both sides by 10: m = 16/10 = 1.6

OpenStudy (anonymous):

thanks